OpenStudy (anonymous):

I need to find the domain and range, the x and y intercepts and the horizontal asymptote, then I need to graph it f(x) =2/x^2-2x-3

jimthompson5910 (jim_thompson5910):

the entire denominator is x^2 - 2x - 3 right?

OpenStudy (anonymous):

Yes!

jimthompson5910 (jim_thompson5910):

ok thanks

jimthompson5910 (jim_thompson5910):

The denominator can't be zero (since you cannot divide by zero), if it were, then... x^2 - 2x - 3 = 0 (x - 3)(x + 1) = 0 x-3 = 0 or x+1 = 0 x = 3 or x = -1

jimthompson5910 (jim_thompson5910):

So the values x = 3 or x = -1 make the denominator zero This means that we must pull these values from the domain

jimthompson5910 (jim_thompson5910):

This also means that the two vertical asymptotes are x = 3 and x = -1

jimthompson5910 (jim_thompson5910):

The horizontal asymptote is y = 0 because the degree of the denominator is larger than the degree of the numerator

jimthompson5910 (jim_thompson5910):

To find the x-intercept, you plug in y = 0 and solve for x To find the y-intercept, you plug in x = 0 and solve for y

OpenStudy (anonymous):

Wouldn't they be the same then?

jimthompson5910 (jim_thompson5910):

no not always

jimthompson5910 (jim_thompson5910):

they're only the same if the graph goes through the origin

OpenStudy (anonymous):

Okay, I'm sorry I don't know how I plug them in, well where, I have to teach myself everything the class is online..

OpenStudy (anonymous):

Do I plug 0 into the original problem?

jimthompson5910 (jim_thompson5910):

there's no teacher? that's stupid

jimthompson5910 (jim_thompson5910):

x-intercept y =2/(x^2-2x-3) 0 =2/(x^2-2x-3) .. plug in y = 0 and solve for x 0(x^2-2x-3) = 2 0 = 2 this is FALSE, so there are no solutions to this particular equation this means that there are no x-intercepts

OpenStudy (anonymous):

It's an online class, It's basically a long page I read and am supposed to understand...

jimthompson5910 (jim_thompson5910):

y-intercept y =2/(x^2-2x-3) y =2/((0)^2-2(0)-3) ... plug in x = 0 and solve for y y =2/(0 - 0 -3) y =2/(-3) y = -2/3 So the y-intercept is (0, -2/3)

jimthompson5910 (jim_thompson5910):

i see

OpenStudy (anonymous):

Okay, thanks:) That very helpful, now I have one more question...How do I find the domain and range?

jimthompson5910 (jim_thompson5910):

well the domain is the set of all real numbers BUT x cannot be 3 and x cannot be -1 remember how these values make the denominator zero, which isn't allowed every other x value works for the domain though

jimthompson5910 (jim_thompson5910):

this is shown above when I found the vertical asymptotes

jimthompson5910 (jim_thompson5910):

the range is the set of all real numbers BUT y cannot be equal to 0 because this is the horizontal asymptote

OpenStudy (anonymous):

Oh, Okay thanks much:)

jimthompson5910 (jim_thompson5910):

yw